OpenStudy (anonymous):

How would I solve log(2x+6) 8 = 3?

OpenStudy (anonymous):

it is \[\large \log_{2x+6}(8)=3\]?

OpenStudy (lucaz):

@GreenEyedGirl ?

OpenStudy (anonymous):

log2x+6(8)=3

OpenStudy (anonymous):

Sorry, I was answering another question. What is the question?

OpenStudy (anonymous):

since i don't know what the question is, i don't know the answer

OpenStudy (anonymous):

I need to solve this logarithmic equation.

OpenStudy (lucaz):

what is the base of the log?

OpenStudy (anonymous):

the binomial inside of the parenthesis

OpenStudy (anonymous):

8 is the argument

OpenStudy (anonymous):

@lucaz ? Still there?

OpenStudy (lucaz):

I don't know how to solve this..

OpenStudy (anonymous):

Do you know someone who would?

OpenStudy (lucaz):

well, maybe you could think of it as (2x+6)^3=8, you want some number that when raised to the 3rd power is equal to 8, 2^3=8, so you want (2x+6)=2

OpenStudy (anonymous):

Can you help me with this one instead? Which of the following is equivalent to log(25) 125? 1.951 2/5 2/3 3/2

OpenStudy (lucaz):